Prev Next

SW&IT Meeting Minutes 2022-06-06

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220530
LR20220530 - Bec, Matthieu 32.0
LR20220530 - Mechtley, Brandon 0.0
LR20220530 - Eychaner, Glenn 32.0
LR20220530 - Edmeades, David 0.0
LR20220530 - Grenz, Paul 0.0
LR20220530 - Hooper, Stephen 0.0
LR20220530 - Zhang, Yang 0.0
LR20220530 - Zhang, Xianyu 32.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • rm507-1 control key (DT for IT)
  • jconsole on obs-x (GE for IT)
  • /scratch on obs4 x2 (OK for IT - DE and LeRoy on top)
  • shark-vis TCS sim / DD overwrite (FP for YZ)
  • bandwidth checker/exerciser (various, MB for IT)
  • lbti-wfs and INDI (MB for MB, SE, PG)
  • what is missing on the new-obs that is available on the old-obs (sciops for SH)
  • dms.tucson outside firewall? 150.135.245.20:53, 80, 443 (MB for IT)

Team reports

MB

  • Last week
    • OC, management, ATT
    • AOS patch
    • asm overcurrent/freq stats
    • OSA nightlog/APU log reduction scripts
  • This week
    • dms refemem recorder @5Hz
    • dms.tucson satellite issue, throttled by cod.edu ?

YZ

  • Back from vacation.
  • Resume some previous TMS work (software and wiki), and perhaps preparing for a test.
  • LBC code study continues.
  • Maybe more discussion with Olga on various LBC/MODS issues we had previously.

PSG

  • Last week
    • Out sick all week.
  • This week
    • Look into PLC code for LBTI - there are still issues that need to be worked out before the run.

GE

  • Last week:
    • IT 8355 (Aladin internal deadlocks), IT 7283 (Subtraction and colormap): I finally figured out the limitations of Aladin threading, and have hopefully been able to resolve these issues permanently. See the ITs for details.
    • IT 5815 (LUCI GUI slowdowns): Dave T. was able to fully demonstrate this issue on obs5 on Friday.
  • This week:
    • I will be taking June 20 and July 21 off, and working half days (mornings) at least 3 days per week between those dates.
    • IT 5815 (LUCI GUI slowdowns)
    • IT 8571, 8529 (Observer Scheduler GUI).

DE

  • Last week:
    • Mountain Veeam proxy and backup
      • Thinking 1x/week, Wednesday daytime
      • vm1.mountain has 4 sockets, taking up 2/3 of our allocation. hv01 has more cores but only 2 sockets, so when we switch to that a primary we'll get 2 licenses back.
    • Petr oopsie, all access revoked
    • Vault issue, caused and solved by Ketiv
    • Obs4 scratch disk, plan to swap it with LD
  • This week:
    • Office trip today to set up JS office hardware
    • LF router
    • Unhappy disk in Synology, will swap it while I'm in
    • Follow up with LD re: obs4
      • Discussion on where to do work like OK is doing ongoing. I was thinking maybe a VM, SH suggested other Obs machines. Will continue when he returns.

XZ

  • Last week
    • DX ASM flat measurements and characterization
    • SX ASM summer activity planning
    • MMC software revision
    • LBTI DX AO support
    • DX ASM bump fix
  • This week
    • MMC SW improment?
    • LBTI DX AO support
    • DX ASM flat characterization
Topic revision: r14 - 06 Jun 2022, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback